OpenStudy (anonymous):

Solve the equation step by step. List each step, along with the property you used during that step. –5 × 1 × 11 × 4.

OpenStudy (anonymous):

you just multiply the numbers.

OpenStudy (dan815):

lol

OpenStudy (dan815):

-220

OpenStudy (anonymous):

step by step

OpenStudy (dan815):

-5*1 = -5 -5*11 = -55 -55*4 = -220